What is 3 times longer than 1 and 5/8
Answer by jim_thompson5910(35256)

1 and 5/8 = 1 + 5/8

1 and 5/8 = 8/8 + 5/8

1 and 5/8 = (8 + 5)/8

1 and 5/8 = 13/8

-------------------------------------------------------

3 times longer than 1 and 5/8 = 3*(1 and 5/8)

3 times longer than 1 and 5/8 = 3*(13/8)

3 times longer than 1 and 5/8 = (3/1)*(13/8)

3 times longer than 1 and 5/8 = (3*13)/(1*8)

3 times longer than 1 and 5/8 = 39/8

Answer as an improper fraction: 39/8

39/8 = (32+7)/8

39/8 = 32/8+7/8

39/8 = 4+7/8

39/8 = 4 and 7/8

Answer as a mixed number: 4 and 7/8